Let X be the number ofcorrect answers if a students guesses randomly from the 5 choices for each of the 25 questions. What is the probability distribution of X?
Answer by stanbon(75887)

Let X be the number ofcorrect answers if a students guesses randomly from the 5 choices for each of the 25 questions. What is the probability distribution of X?
----
It's a binomial distribution with
u = np = 25*(1/5) = 5
and
s = sqrt(npq) = sqrt(5(4/5)) = 2
------------------------------------
Cheers,
Stan H.
========================================